OpenStudy (anonymous):

the line through (8,-2) perpendicular to the line y=7-2x

OpenStudy (anonymous):

Perpendicular means it has a negative reciprocal of the slope, which means the slope with be 1/2

OpenStudy (anonymous):

Now since u have the slope, you would plug it into the point-slope formula

OpenStudy (anonymous):

Y-(-2)=1/2(X-8) Y+2=1/2x-4 y=1/2x-6

OpenStudy (anonymous):

Their is your line